Find the common difference, the fourth term, and the "n"th term in the arithmetic sequence 80,76,72,....
Answer by richwmiller(17219)

An = A1 + (n - 1)d
d=-4
An = 80-4*( n - 1)
An=80-4n+1
An=81-4n